OpenStudy (anonymous):

If 4f(x)+f(3-x)=x^2 what is f(x)?

OpenStudy (anonymous):

I think you solve by switching the x to (3-x) and the 3-x to x if that makes sense. I don't know where to go from there.

OpenStudy (anonymous):

What level math are you doing?

OpenStudy (anonymous):

College Algebra.

OpenStudy (amistre64):

it looks to be that x = 3-x if anything, its worth a shot :) 2x = 3; x = 3/2 4(3/2) + 3/2 ?= 9/4 12/2 + 3/2 = 15/2 ..... well that dint work out lol

OpenStudy (amistre64):

its hard to tell without more information ... at least for me it is

OpenStudy (anonymous):

I have noooooo idea. Lol. I googled and found a similar problem. It might have something to do with synthetic division

OpenStudy (anonymous):

Please share!

OpenStudy (amistre64):

i cant make sense of it at the moment :)

OpenStudy (anonymous):

well.. i don't know the url. I googled it the other day =T but this might help. http://au.answers.yahoo.com/question/index?qid=20110522002904AATkbUG

OpenStudy (anonymous):

I had this problem the other day for an assignment, 5f(x)+f(3-x)=x^2 and the right answer was supposed to come out to f(x)=1/6x^2+1/4x-3/8 .

OpenStudy (anonymous):

I'm just working the example problem trying to figure out how they got their answer. If I figure it out I'll give you a yell!

OpenStudy (anonymous):

PLEASE DO! lol

OpenStudy (anonymous):

I ran out of time on the assignment :/ But the answer is supposed to be 1/5x^2 + 2/5x - 3/5.

OpenStudy (anonymous):

do you have any of the steps? think this has to do with finding the inverse yes?

OpenStudy (anonymous):

or are you supposed to know that \[f(x)=ax^2+bx+c\] and solve for a, b and c?

OpenStudy (anonymous):

Unfortunately, no. It wasn't a problem that was solved in class.

OpenStudy (anonymous):

ok was the first term \[\frac{1}{5}x^2\]?

OpenStudy (anonymous):

From the example I found online it looks like a combination of that formula above f(x)=ax^2+bx+c and synthetic division.

OpenStudy (anonymous):

if that is the right first term we can do this

OpenStudy (anonymous):

It was!

OpenStudy (anonymous):

whew. ok ready? here we go!

OpenStudy (anonymous):

i guess you have to know that \[f(x)\] is going to be a quadratic of some sort. that is \[f(x)=ax^2+bx+c\]

OpenStudy (anonymous):

we just don't know what a, b and c are and that is what we have to find.

OpenStudy (anonymous):

so we compute \[4f(x)=4ax^2+4bx=4c\] easy enough. now what is \[f(3-x)\]? \[f(3-x)=a(3-x)^2+b(3-x)+c\] and so \[4f(x)+f(3-x)=4ax^2+4bx+c+a(3-x)^2+b(3-x)+c\]

OpenStudy (anonymous):

sorry typo on first line. it should be \[4f(x)=4ax^2+4bx+4c\]

OpenStudy (anonymous):

and we know that all that muck is equal to \[x^2\] because that is what we are told

OpenStudy (anonymous):

now the fun job is to multiply out. you get \[4ax^2+4bx+4c+9a-6ax+ax^2 +3b-bx+c=x^2\]

OpenStudy (anonymous):

now we collect like terms. for the \[x^2\] term we have \[4ax^2+ax^2=5ax^2\] and we know that this is equal to \[x^2\] which means \[5a=1\] so \[a=\frac{1}{5}\]

OpenStudy (anonymous):

did i lose you yet?

OpenStudy (anonymous):

I'm following you, but why is c not c(3-x) as well?

OpenStudy (anonymous):

oh because c is a constant. if i replace x by 3-x is \[ax^2+bx+c\] i get \[a(3-x)^2+b(3-x)+c\]

OpenStudy (anonymous):

hope it is more or less clear what i did in the second part. i was not multiplying by 3-x i was replacing x by 3-x

OpenStudy (anonymous):

Oh. That makes more sense. The first part was multiplication and the second part was filling in for x then?

OpenStudy (anonymous):

yes the first part was multiply by 4 because it was \[4f(x)\] the second part was replacing x by 3-x because it was \[f(3-x)\]

OpenStudy (anonymous):

so now at least we know that \[a=\frac{1}{5}\] by equating coefficients. you have on one side \[5ax^2\] and on the other side you have \[x^2\] so as i said \[5a=1\] \[a=\frac{1}{5}\]

OpenStudy (anonymous):

now we do the same for the "x" term. there are no x's on the right because you just have \[x^2\]

OpenStudy (anonymous):

on the left i have to scroll up and see what we get

OpenStudy (anonymous):

yes on the left we have for the "x" term \[4bx-6ax-bx\] \[3bx-6ax\] \[(3b-6a)x\]

OpenStudy (anonymous):

and we know that \[3b-6a=0\] because we should have no x term

OpenStudy (anonymous):

\[3b=6a\] \[3b=\frac{6}{5}\] \[b=\frac{2}{5}\] this because we already knew that \[a=\frac{1}{5}\]

OpenStudy (anonymous):

is that what we should have for b?

OpenStudy (anonymous):

and finally we do the same thing with the constant. the constant on the right is 0. the constant on the left is ... let me scroll up.

OpenStudy (anonymous):

It's right. I'm a little bit stuck on how you figured out that 4ax^2 +ax^2=5ax^2

OpenStudy (anonymous):

oh i can explain that. we call that "combining like terms" yes?

OpenStudy (anonymous):

and the 3b-6a part.

OpenStudy (anonymous):

4 somethings and 1 something is 5 somethings.

OpenStudy (anonymous):

I get that.

OpenStudy (anonymous):

yes? \[4x+x=5x\] \[4ax+ax=5ax\] \[4ax^2+ax^2=5ax^2\] tex

OpenStudy (anonymous):

k?

OpenStudy (anonymous):

now lets go all the way back to \[4ax^2+4bx+4c+9a−6ax+ax^2+3b−bx+c=x^2\]

OpenStudy (anonymous):

and rewrite the left and side in standard form. in other words combine like terms

OpenStudy (anonymous):

Ohh. Okay I think I get it. I was confused about what the 9a-6ax+ax^2+3b-bx+c what about.

OpenStudy (anonymous):

\[4ax^2+ax^2+4bx-6ax-bx+4c+9a+3b+c=x^2\]

OpenStudy (anonymous):

\[5ax^2+(3-6a)x+(5c+9a+3b)=x^2\]

OpenStudy (anonymous):

damned typo!

OpenStudy (anonymous):

\[5ax^2+(3b-6a)x+(5c+9a+3b)=x^2\]

OpenStudy (anonymous):

this tells us that \[5a=1, 3b-6a=0,5c+9a+3b=0\]

OpenStudy (anonymous):

is that step clear?

OpenStudy (anonymous):

because thats is what you get when you combine like terms on the left, and on the right you have \[1x^2+0x+0\]

OpenStudy (anonymous):

so if that part is ok it is gravy from here on in. first solve for a, then solve for b, then solve for c

OpenStudy (anonymous):

as you can see setting this up is a pain. solving not so bad

OpenStudy (anonymous):

wait, why exactly is that 0? :/

OpenStudy (anonymous):

ok you are given that \[4f(x)+f(3-x)=x^2\] yes?

OpenStudy (anonymous):

and we just rewrote the left hand side but the right hand side is still just \[x^2\]

OpenStudy (anonymous):

so whatever "x" terms we have on the left, on the right there are none

OpenStudy (anonymous):

on the left side we collect like terms and get \[(3a-6b)x\] but on the right side we have \[0x\]

OpenStudy (anonymous):

no x term on the right, \[3a-6b\]of them on the left.

OpenStudy (anonymous):

that tells us that \[3a-6b=0\]

OpenStudy (anonymous):

analogous to how we knew that \[5a=1\] on the left we have \[5ax^2\] on the right we have \[x^2\] so we know \[5a=1\]

OpenStudy (anonymous):

on the left we have \[(3a-6b)x\] on the right we have \[0x\] on the left the "constant" is \[9a+3b+5c\] on the right the "constant" is 0

OpenStudy (anonymous):

that's pretty clear now

OpenStudy (anonymous):

ok good. so we so\ lve one at a time. first we found \[a=\frac{1}{5}\] then we know that \[3b-6a=0\] \[3b=6a\] \[3b=\frac{6}{5}\] \[b=\frac{2}{5}\]

OpenStudy (anonymous):

and finally we solve \[5c+9a+3b=0\] and now this is too much arithmetic for me but the idea is that we know this is \[5c+\frac{9}{5}+\frac{6}{5}=0\] etc

OpenStudy (anonymous):

might as well finish... \[5c+\frac{15}{5}=0\] \[5c+3=0\] \[5c=-3\] \[c=-\frac{3}{5}\]

OpenStudy (anonymous):

I got it. :D

OpenStudy (anonymous):

Thank you so much!